OpenStudy (anonymous):

find the difference 1/8(-8/3-1/8)?

OpenStudy (pfenn1):

Is this the expression?\[\frac18 \times \left( -\frac83-\frac18 \right)\]

OpenStudy (anonymous):

no instead of multiplying I need to subtract the rest is correct =)

OpenStudy (pfenn1):

So you mean...\[\frac18 -\frac83-\frac18\]

OpenStudy (anonymous):

yes, but with the brackets, does it make a difference?

OpenStudy (pfenn1):

It shouldn't. If the brackets were like this it would...\[\frac18-\left( \frac83-\frac18 \right)\]But you could leave the brackets as find the common denominator of 8/3 and 1/8. You would get the same answer.

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (anonymous):

okay thanks for clearing that out for me =)

OpenStudy (pfenn1):

np

OpenStudy (anonymous):

What would be the correct answer?

OpenStudy (pfenn1):

-8/3